OpenStudy (anonymous):

how can i isolate c when a=b(1/c-1/d)

jhonyy9 (jhonyy9):

a = b/c -b/d a + b/d = b/c (ad +b)/d =b/c c= bd/(ad+b)

OpenStudy (anonymous):

thank you sooooo much :)

jhonyy9 (jhonyy9):

was my pleasure